proof the expression below : MSE (0) = Var (ôh + b (62
Answer #1

since MSE is given by

MSE(\hat{\theta})=E[(\hat{\theta}-\theta)^2] =E[(\hat{\theta}-\theta-E(\hat{\theta})+E(\hat{\theta}))^2]=E[((\hat{\theta}-E(\hat{\theta}))+(E(\hat{\theta})-\theta ))^2]

  =E[((\hat{\theta}-E(\hat{\theta}))^2]+E[(E(\hat{\theta})-\theta ))^2]+2E[(\hat{\theta}-E(\hat{\theta})*(E(\hat{\theta})-\theta ))]

  =E[((\hat{\theta}-E(\hat{\theta}))^2]+E[(E(\hat{\theta})-\theta ))^2]+0 ,.............(as..(E(\hat{\theta})-\theta )=const,E(\hat{\theta}-E(\hat{\theta}))=0)

=E[((\hat{\theta}-E(\hat{\theta}))^2]+E[(E(\hat{\theta})-\theta ))^2]

=Var(\hat{\theta}) +(bias(\hat{\theta}))^2

Where

Var(\hat{\theta})=E[(\hat{\theta}-E(\hat{\theta}))^2],Bias=E(\hat{\theta})-\theta